### v1

# A Pioneering Study and An Innovative Information Theory-based Approach to Enhance The Transparency in Phishing Detection

Abstract.
Phishing attacks have become a serious and challenging issue for detection, explanation, and defense. Despite more than a decade of research on phishing, encompassing both technical and non-technical remedies, phishing continues to be a serious problem. Nowadays, AI-based phishing detection stands out as one of the most effective solutions for defending against phishing attacks by providing vulnerability (i.e., phishing or benign) predictions for the data. However, it lacks explainability in terms of providing comprehensive interpretations for the predictions, such as identifying the specific information that causes the data to be classified as phishing. To this end, we propose an innovative deep learning-based approach for email (the most common phishing way) phishing attack localization. Our method can not only predict the vulnerability of the email data but also automatically figure out and highlight the most important and phishing-relevant information (i.e., sentences) in each phishing email. The selected information indicates useful explanations for the vulnerability of the phishing email data. The rigorous experiments on seven real-world email datasets show the effectiveness and advancement of our proposed method in providing comprehensive explanations (by successfully figuring out the most important and phishing-relevant information in phishing emails) for the vulnerability of corresponding phishing data with higher performances from nearly (1% to 3%) and (1% to 4%) in two main Label-Accuracy and Cognitive-True-Positive measures, respectively, compared to the state-of-the-art potential baselines.
1. Introduction
Phishing attacks (i.e., attempts to deceitfully get personal and financial information such as usernames, passwords, and bank accounts through electronic communication with malicious intentions) have become a serious issue. Nowadays, there are various ways to conduct phishing attacks; the most common method is through the use of emails. Email phishing is crafted to trigger psychological reactions in the users by using persuasion techniques via cognitive principles  (Van Der Heijden and Allodi, 2019) such as scarcity, consistency, and authority.
According to recent reports, there has been a notable rise in the occurrence of increasingly sophisticated phishing attacks, presenting more formidable challenges for detection and defense. The FBI Internet Crime Report  (Federal Bureau of Investigation, 2022) reveals that phishing attacks were the most common cyber-crime type in the United States causing substantial economic losses every year from 2018. Similarly, the Anti-phishing Working Group (APWG) (Anti-Phishing Working Group, 2022) reports that the number of phishing attacks has grown by more than 150% per year since the beginning of 2019. The year 2022 was a record for phishing, with the APWG logging more than 4.7 million attacks. These reports highlight the danger of phishing attacks and the need for a greater understanding of this serious problem.
The widespread adoption of artificial intelligence (i.e., using machine learning-based and deep learning-based approaches) has brought substantial influences and great success in various domain applications such as autonomous driving (Chen et al., 2023b)), data generations (Devlin et al., 2018; Radford et al., 2019; Raffel et al., 2019)), drug discovery (Paul et al., 2021)), and malware vulnerability detection (Li et al., 2018; Nguyen et al., 2021b)). By leveraging the power of machine learning and deep learning, there have been many efforts proposed for solving phishing attack problems from phishing attack detection (Ramanathan and Wechsler, 2013; Le et al., 2018; Tyagi et al., 2018; Sahingoz et al., 2019; Yang et al., 2019; Das et al., 2019; Xiao et al., 2020; Yang et al., 2021) to the investigation of the cognitive bias’s impact used in email phishing (Parsons et al., 2015; Van Der Heijden and Allodi, 2019). It has been proven that using machine learning and deep learning-based algorithms to detect phishing attacks in the early stages is one of the most effective solutions for preventing and reducing the negative effects caused. Recently, with ongoing capabilities and improvements in Natural language processing (NLP) and the popularity of large language models, e.g., GPT2 (Radford et al., 2019), BERT (Devlin et al., 2018), and T5 (Raffel et al., 2019), some of the early attempts use large language models for phishing attack-related issues have been introduced, e.g., scam-baiting mail servers that can conduct scam-baiting activities automatically (Chen et al., 2023a).
Despite more than a decade of research on phishing (i.e., encompassing both technical and non-technical remedies), it still continues to be a serious problem: (i) the problem itself may be intractable, the technical approaches so far may have missed important characteristics (e.g., psychological manipulation techniques used by attackers) of the problem, and (ii) phishing exploits the human as the weakest link so purely technical approaches may not be sufficient (Das et al., 2019). To overcome these problems, in addition to enhancing human understanding of phishing, phishing attack detection appears as one of the most effective solutions to help human users defend against phishing attacks by providing the vulnerability prediction (i.e., phishing or benign) before they take any actions following the content mentioned in the corresponding data (e.g., the email’s content). Although phishing detection methods can predict the data’s vulnerability label (i.e., phishing or benign), these predictions often lack the explainability to offer meaningful interpretations (i.e., which information causing the data phishing) to the users. Motivated by this problem, in this paper, we study the following research question:
”In addition to predicting the vulnerability (i.e., phishing or benign) of the data, how to derive an effective deep learning-based method that can also automatically figure out and highlight the important and phishing information in each phishing data for providing a useful and comprehensive explanation about the vulnerability of the data to users? (we name this problem as phishing attack localization)”
To this end, in the scope of our paper, we study the phishing attack localization problem on phishing emails (i.e., the most common phishing way) where we propose an innovative information theory-based model to solve the problem. Our method not only can detect the vulnerability (i.e., phishing or benign) of the email but also automatically learns and identifies the most important and phishing-relevant information (e.g., sentences) in each phishing email. The selected phishing-relevant information (e.g., sentences) helps provide useful explanations about the vulnerability of the phishing email data. It is worth noting that the ability to figure out the important and phishing-relevant information that causes the associated email data to be classified as phishing to provide a corresponding comprehensive interpretation is the main difference between phishing attack localization and phishing attack detection.
In summary, our key contributions are as follows:
  • We study an important problem of phishing attack localization aiming to tackle and improve the explainability (transparency) of phishing detection. Automated machine learning and deep learning-based techniques for this problem have not yet been well studied.
  • We propose a novel deep learning-based framework derived from an information-theoretic perspective and information bottle-neck theory for phishing attack localization. Our proposed approach can work effectively in a weakly supervised setting (refer to Section 3.1 for details), hence providing an important practical solution for defeating phishing attacks.
  • Besides proposing a deep learning-based approach for phishing attack localization, based on the explainable machine learning and email phishing domain knowledge, we propose some appropriate measures including Label-Accuracy, Cognitive-True-Positive, and SAC (refer to Section 4.1.3 for details) for the problem.
  • We comprehensively evaluate our method on real-world phishing email datasets. The rigorous and extensive experiments show the superiority of our method over the state-of-the-art baselines.
2. Related Work
2.1. Phishing attack detection
Machine learning has emerged as a natural choice for addressing classification challenges, including the detection of phishing attempts. Over the past decades, extensive research has been dedicated to this area. Numerous machine learning methods have been employed to tackle the issue, creating models capable of classifying content as either benign or phishing (Tyagi et al., 2018; Rao and Pais, 2019; Li et al., 2019; Zamir et al., 2020). These models leverage classification principles and carefully choose features that represent the content’s distinctive characteristics. However, machine learning-based methods are heuristics and require the knowledge of domain experts to manually engineer features from the data that can be outdated and biased. To overcome these limitations, deep learning-based solutions have been proposed for the phishing attack detection problem (Xiao et al., 2020; Abdelnabi et al., 2020; Yang et al., 2019; Lin et al., 2021) and have shown significant advantages over machine learning approaches that use hand-crafted features. AI-based phishing detection methods can predict the vulnerability label (i.e., phishing or benign) of the data (e.g., emails). However, they lack the ability to provide comprehensive explanations for their predictions.
2.2. Phishing attack localization
Automated deep learning-based techniques for the phishing attack localization problem (i.e., to tackle and improve the explainability (transparency) of phishing detection) have not yet been well studied. The interpretable machine-learning research appears to be an appropriate direction for dealing with the phishing attack localization problem. In short, interpreting approaches (e.g., (Caruana et al., 2015; Rich, 2016; Marco T. Ribeiro, 2016; Chen et al., 2018a; Bang et al., 2021a; Yoon et al., 2019b; Nguyen et al., 2021a; Jethani et al., 2021a)) are used for explaining the behavior of deep learning-based systems or the ground truth label of the data by automatically learning and figuring out the most important information (e.g., attributions or features) existing from the data that are responsible in causing the corresponding decision of black-box models and the ground truth label. For example, in sentiment analysis, interpretable machine-learning approaches (e.g., (Chen et al., 2018a; Vo et al., 2023)) help to give a comprehensive explanation for the review (positive or negative) of a movie by figuring out and highlighting the top important keywords or sentences. In our paper, via addressing phishing attack localization, we not only aim to predict the vulnerability (i.e., phishing or benign) of the emails but also aim to provide a comprehensive interpretation by figuring out and highlighting the most crucial phishing-relevant information causing the emails to be classified as phishing.
In our paper, we demonstrate that interpretable machine learning methods are strongly suitable for phishing attack localization in automatically learning and figuring out the most important features of the data obtained from the models to explain the model’s prediction decision. We compared the performance of our proposed method with some popular and state-of-the-art interpretable machine learning approaches including L2X ((Chen et al., 2018b)), INVASE ((Yoon et al., 2019a), ICVH ((Nguyen et al., 2021a), VIBI ((Bang et al., 2021b)), and AIM ((Vo et al., 2023)) (please refer to Section 4.1.4 in the experiments section for our brief description of these methods).
3. The proposed approach
3.1. The problem statement
We denote an email data sample as $X=[\mathbf{x}_{1},...\mathbf{x}_{L}]$ consisting of $L$ sentences. In the scope of our paper, we consider each email as a sequence of sentences. We assume that $X^{\prime}$s vulnerability label $Y\in\{0,1\}$ (i.e., $1$: phishing and 00: benign). In the context of phishing attack localization, we aim to build an automatic AI-based approach mainly not only for detecting the vulnerability $Y$ of the email $X$ but also for automatically learning and figuring out the important and phishing-relevant information (i.e., sentences) denoted by $\tilde{X}$ (a subset of $X$) causing $X$ phishing.
It is worth noting that for almost all publicly available phishing-relevant data (e.g., emails), there are only labels related to the data’s vulnerability (phishing or benign) by domain experts with the assistance of machine learning or deep learning tools. We almost do not have the ground truth of phishing information (the information truly causes the data phishing). In the phishing attack localization problem, we name this context as a weakly supervised setting where during the training process, we only use the vulnerability label of the data while not using the ground truth of phishing information existing in the data when solving the phishing attack localization problem.
3.2. Methodology
We name our proposed method as AI2TALE. In the following, we present details of how our AI2TALE  method works and addresses the phishing attack localization problem to tackle and improve the explainability (transparency) of phishing detection. An overall visualization of our method is depicted in Figure 1.
Figure 1. A visualization of our proposed AI2TALE  method for the phishing attack localization problem.
3.2.1. Learning to select important and phishing-relevant information and the training principle
Phishing-relevant information selection process
As shown in Figure 1, the first part of our method is the selection network $\zeta$. It aims to learn and figure out the most important and label-relevant information (i.e., sentences) in each email in an automatic and trainable manner. Note that, in terms of the phishing email, the key selected information stands for the phishing-relevant information causing the email phishing.
Giving an email $X$ consisting of $L$ sentences from $\mathbf{x}_{1}$ to $\mathbf{x}_{L}$ (i.e., in practice, each sentence $\mathbf{x}_{i}$ is represented as a vector using a learnable embedding method, please refer to the data processing and embedding section 4.1.2 for details), to figure out the important and label-relevant sentences $\tilde{X}$ in $X$, we introduce a selection process $\zeta$ (i.e., it is learnable and maps $\mathbf{R}^{L}\mapsto[0,1]^{L}$) aiming to learn a set of independent Bernoulli latent variables $\mathbf{z}\in\{0,1\}^{L}$ representing the importance of the sentences to the email’s vulnerability $Y$. Specifically, each element $z_{i}$ in $\mathbf{z}=\{z_{i}\}_{i=1}^{L}$ indicates whether $\mathbf{x}_{i}$ is related to the vulnerability $Y$ of $X$ (i.e., if $z_{i}$ is equal to $1$, the sentence $\mathbf{x}_{i}$ plays an important role causing the vulnerability $Y$).
To construct $\mathbf{z}=\{z_{i}\}_{i=1}^{L}$, we model $\mathbf{z}\sim\mathrm{MultiBernoulli}(\mathbf{p})=\prod_{i=1}^{L}\text{Bernoulli}(p_{i})$, which indicates $\boldsymbol{x}_{i}$ is related to the vulnerability $Y$ of $X$ with probability $p_{i}$ where $p_{i}=\omega_{i}\left(X;\alpha\right)$ with $\omega$ is a neural network parameterized by parameter $\alpha$. It is worth noting that the neural network $\omega$ takes $X$ as input and outputs corresponding $\mathbf{p}=\{p_{i}\}_{i=1}^{L}$. With $\mathbf{z}$, we construct $\tilde{X}=\zeta\left(X\right)$ (i.e., the subset sentences that lead to the vulnerability $Y$ of the function $X$) by $\tilde{X}=X\odot\mathbf{z}$, where $\odot$ represents the element-wise product. Note that we can view the selector model as a distribution $q_{sel}(\mathbf{z}|X;\alpha)$ over a selector variable $\mathbf{z}$, which indicates the important features for a given sample of $X$.
Reparameterization for continuous optimization
Recall that the selected sentences $\tilde{X}$ are constructed by $\tilde{X}=\mathbf{z}\odot X$ where $\mathbf{z}\sim\mathrm{MultiBernoulli}(\mathbf{p})$ with $\mathbf{p}$ is the output of the selection process $\zeta$ parameterised by a neural network $\omega(.,\alpha)$.
To make this computational process (i.e., the process consists of sampling operations from a Multi-Bernoulli distribution) continuous and differentiable during training, we apply the temperature-dependent Gumbel-Softmax trick (Eric Jang, 2016; Maddison et al., 2016) for relaxing each Bernoulli variable $z_{i}$. We sample $z_{i}\left(X;\alpha\right)\sim\text{Concrete}(\omega_{i}(X;\alpha),1-\omega_{i}(X;\alpha))$:
$z_{i}\left(X;\alpha\right)=\frac{\exp\{(\log\omega_{i}+a_{i})/\tau\}}{\exp\{(\log\omega_{i}+a_{i})/\tau\}+\exp\{(\log\left(1-\omega_{i}\right)+b_{i})/\tau\}}$
where we denote $\omega_{i}\left(X;\alpha\right)$ as $\omega_{i}$ while $\tau$ is a temperature parameter (i.e., that allows us to control how closely a continuous representation from a Gumbel-Softmax distribution approximates this from the corresponding discrete representation from a discrete distribution (e.g., the Bernoulli distribution)), random noises $a_{i}$ and $b_{i}$ independently drawn from Gumbel distribution $G=-\log(-\log u)$ with $\ u\sim\textbf{Uniform}(0,1)$.
Mutual information for guiding the selection process
In information theory (Shannon, 1998; Cover and Thomas, 2006), mutual information is used to measure the mutual dependence between two random variables. In particular, it quantifies the information obtained about one random variable by observing the other. To illustrate, consider a scenario where A denotes the outcome of rolling a standard 6-sided die, and B represents whether the roll results in an even number (0 for even, 1 for odd). Evidently, the information conveyed by B provides insights into the value of A, and vice versa. In other words, these random variables possess mutual information.
Leveraging this property of mutual information and inspired by (Chen et al., 2018c; Nguyen et al., 2021a), we maximize the mutual information between $\tilde{X}$ and $Y$ as mentioned in Eq. (1) with the intuition is that by using the information from $Y$, the selection process $\zeta$ will be learned and enforced to obtain the most meaningful $\tilde{X}$ (i.e., $\tilde{X}$ can predict the vulnerability $Y$ of $X$ correctly).
By viewing $\tilde{X}$ and $Y$ as random variables, the selection process (model) $\zeta$ is learned by maximizing the mutual information between $\tilde{X}$ and $Y$ as follows:
(1) $\max_{\zeta}\,\mathbf{I}(\tilde{X},Y)$
Following (Cover and Thomas, 2006), we expand Eq. (1) further as the Kullback-Leibler divergence (i.e., it measures the relative entropy or difference in information represented by two distributions) of the product of marginal distributions of $\tilde{X}$ and $Y$ from their joint distribution:
$\mathbf{I}(\tilde{X},Y)=$ $\int p(\tilde{X},Y)\log\frac{p(\tilde{X},Y)}{p(\tilde{X})p(Y)}d\tilde{X}dY$
(2) $\geq$ $\int p(Y,\tilde{X})\log\frac{q(Y\mid\tilde{X})}{p(Y)}dYd\tilde{X}$
In practice, estimating mutual information is challenging as we typically only have access to samples but not the underlying distributions. Therefore, in the above derivation, we apply to use a variational distribution $q(Y|\tilde{X})$ to approximate the posterior $p(Y\mid\tilde{X})$, hence deriving a variational lower bound of $\mathbf{I}(\tilde{X},Y)$ for which the equality holds if $q(Y\mid\tilde{X})=p(Y\mid\tilde{X})$. This can be further expanded as:
$\mathbf{I}(\tilde{X},Y)\geq$ $\int p(Y,\tilde{X},X)\log\frac{q(Y\mid\tilde{X})}{p(Y)}dYd\tilde{X}dX$
(3) $=$ $\mathbf{E}_{X,Y}\mathbf{E}_{\tilde{X}|X}[\log q(Y|\tilde{X})]+\text{const}$
To model the conditional variational distribution $q(Y|\tilde{X})$, we introduce a classifier implemented with a neural network $f(\tilde{X};\beta)$, which takes $\tilde{X}$ as input and outputs its corresponding label $Y$ (i.e., we view the classifier model as a distribution $q_{class}(Y|\tilde{X};\beta)$). Our objective is to learn the selection process (model) as well as the classifier to maximize the following objective function:
(4) $\underset{\alpha,\beta}{\text{max}}\left\{\mathbf{E}_{X,Y}\mathbf{E}_{\mathbf{z}\sim q_{sel}(\mathbf{z}|X;\alpha)}[\text{log}q_{class}(Y|X\odot\mathbf{z};\beta)]\right\}$
It is important to highlight that mutual information plays a crucial role in fostering collaboration between the classifier and the selection process during training. Specifically, the classifier focuses on identifying a subset of features crucial for assigning labels to data samples while the selection process is tailored to choose the optimal subset based on the feedback provided by the classifier.
3.2.2. Benefits as well as potential weaknesses of the mutual information training principle and our innovative solutions
With the training principle mentioned in Eq. (1), we aim to maximize the mutual information between $\tilde{X}$ and $Y$ for guiding the whole training process to figure out the sentences related to an email’s vulnerability. Recall that during this training process, the classifier learns to identify a subset of sentences leading to an email sample’s vulnerability label while the selection process is designed to select the best subset according to the feedback of the classifier.
In short, this joint training process between the classifier $f(.,\beta)$ and the selection network $\omega(.,\alpha)$ brings benefits for selecting the important and phishing-relevant sentences from phishing email; However, we observe two potential limitations of this training principle as follows:
Obtaining a superset of phishing-relevant sentences
The first limitation of the training principle mentioned in Eq. (1) is that it does not theoretically guarantee to eliminate sentences unrelated to the vulnerability of a specific email. Therefore, the set of selected sentences can be a superset of the true phishing-relevant sentences. In the worst case, a selection process can always select all the sentences in an email, which is still a valid solution for the above maximization mentioned in Eq. (1). To deal with this problem to let the model be able to successfully select and highlight true phishing-relevant information, inspired by (Tishby et al., 2000; Slonim and Tishby, 2000) about using information bottleneck theory, we propose an additional term for training the selection process (model) $\zeta$, derived from the following principle:
(5) $\max_{\zeta}\,(\mathbf{I}(\tilde{X},Y)-\lambda\mathbf{I}(X,\tilde{X}))$
where $\lambda$ is a hyper-parameter indicating the weight of the second mutual information.
By minimizing the mutual information between $X$ and $\tilde{{X}}$, we encourages $\tilde{{X}}$ to be as ”different” to $X$ as possible. In other words, the selection process prefers to select a smaller subset that excludes the sentences unrelated to vulnerability $Y$ of the corresponding data $X$. Accordingly, we can derive an upper bound of the minimization between mutual information between $X$ and $\tilde{X}$:
$\mathbf{I}(\tilde{X},X)=\int p(\tilde{X},X)\log\frac{p(\tilde{X}|X)}{p(\tilde{X})}d\tilde{X}dX$
(6) $\leq\mathbf{E}_{X}\mathbf{E}_{\tilde{X}|X}[\log\frac{p(\tilde{X}|X)}{r(\tilde{X})}]$
for any distribution $r(\tilde{X})$.
We then further derive $\mathbf{I}(\tilde{X},X)$ as the Kullback-Leibler divergence of the product of marginal distributions of $\tilde{X}$ and $X$ from their joint distribution:
$\mathbf{E}_{X}\mathbf{E}_{\tilde{X}|X}[\log\frac{p(\tilde{X}|X)}{r(\tilde{X})}]=\mathbf{E}_{X}\mathbf{E}_{\tilde{X}|X}[\sum_{i=1}^{L}\log\frac{p(\tilde{\boldsymbol{x}_{i}}|X)}{r(\tilde{\boldsymbol{x}_{i}})}]$
(7) $\,\,\,\,\,\,\,\,\,\,\,\,\,\,=\sum_{i=1}^{L}\mathbf{E}_{X}[D_{KL}(p(\tilde{\boldsymbol{x}_{i}}|X)\|r(\tilde{\boldsymbol{x}_{i}}))]$
Minimizing $\mathbf{I}(\tilde{X},X)$ is now equivalent to minimizing the KL divergence between $p(\tilde{\boldsymbol{x}_{i}}|X)$ and $r(\tilde{\boldsymbol{x}_{i}})$. Therefore, one can view $r(\tilde{\boldsymbol{x}_{i}})$ as the prior distribution, which is constructed by $r(\tilde{\boldsymbol{x}_{i}})=\mathcal{N}(\tilde{\boldsymbol{x}}_{i}|0,\sigma^{2})$. Given the fact that $p(\tilde{\boldsymbol{x}}_{i}|X)$ is a Gaussian mixture distribution (i.e., between $p_{i}\mathcal{N}(\tilde{\boldsymbol{x}}_{i}\mid\boldsymbol{x}_{i},\sigma^{2})$ and ($1-p_{i})\mathcal{N}(\tilde{\boldsymbol{x}_{i}}\mid 0,\sigma^{2})$ where $\sigma>0$ is a small number), the intuition is that the prior prefers the small values centered at 00. In this way, $p(\tilde{\boldsymbol{x}}_{i}|X)$ is encouraged to select fewer sentences. Moreover, the KL divergence $D_{KL}(p(\tilde{\boldsymbol{x}}_{i}|X)\|r(\tilde{\boldsymbol{x}}_{i}))$ can be computed by the following approximation (Gal and Ghahramani, 2016):
(8) $\frac{\omega_{i}(X;\alpha)}{2\sigma^{2}}\left\|\boldsymbol{x}_{i}^{2}\right\|+(\log\,\sigma+\frac{1}{2}\sigma^{2})+\text{const}$
Combining $\mathbf{I}(\tilde{X},Y)$ and $\mathbf{I}(X,\tilde{X})$ by $\max\,(\mathbf{I}(\tilde{X},Y)-\lambda\mathbf{I}(X,\tilde{X}))$, we get a unified training objective:
$\max_{\alpha,\beta}\{\mathbf{E}_{X,Y}\mathbf{E}_{\mathbf{z}\sim q_{sel}(\mathbf{z}|X;\alpha)}[\text{log}q_{class}(Y|X\odot\mathbf{z};\beta)]$
(9) $\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,-\lambda\mathbf{E}_{X}[\sum_{i=1}^{L}\frac{\omega_{i}\left(X;\alpha\right)}{2\sigma^{2}}\left\|\boldsymbol{x}_{i}^{2}\right\|\}$
Encoding the vulnerability label via its selections instead of via truly meaningful information
The selected features obtained from the joint training process of the classifier $f(.,\beta)$ and the selection network $\omega(.,\alpha)$ cause the other potential limitation for the training principle mentioned in Eq. (1). In particular, the predictions of the classifier $f(.,\beta)$ can be based more on the features selected from the selection network $\omega(.,\alpha)$ than the underlying information contained in the features. In this case, the selected information (i.e., sentences) can be any subsets of the entire sentences and can be less likely to be meaningful ones from the data.
To deal with this problem and ensure the learnable selection process respecting the data distribution to select the meaningful and label-relevant information (e.g., sentences) of the data, in addition to learning the classifier jointly with the selection network as mentioned in Eq. (9), inspired by (Jethani et al., 2021b), we propose to learn the classifier model $f(.,\beta)$ disjointly to approximate the ground truth conditional distribution of $Y$ given $X_{R}$ where $X_{R}=X\odot\mathbf{r}$ with $\mathbf{r}\sim\text{MultiBernoulli}(0.5)$ (denoted by $\mathbf{r}\sim\text{B}(0.5)$ for short). This procedure helps adjust the classifier to let it not only be affected by the information obtained from the selection network but also based on the information from the data to update its parameters. That helps prevent the problem of encoding the vulnerability label via its selections to improve the data representation learning process. We name this procedure as a data-distribution mechanism. This procedure is then expressed as learning $q_{class}(.;\beta)$ to maximize:
(10) $\mathbf{E}_{X,Y}\mathbf{E}_{\mathbf{r}\sim\text{B}(0.5)}\{\text{log}q_{class}(Y|X\odot\mathbf{r};\beta)\}$
[…]
3.2.3. A summary of our AI2TALE  method
Algorithm 1 shows the details of our proposed AI2TALE  method in the training phase. It is worth noting that during the training process, our model is trained to learn and figure out the most important and phishing-relevant sentences in each corresponding email without using any information about the ground truth of phishing-relevant sentences. This shows a great advantage of using our method for phishing attack localization in practice because in reality, to almost all publicly available phishing-relevant email datasets, there is only information about the vulnerability $Y$ (i.e., phishing or benign) of the data by domain experts with the help of machine learning and deep learning tools.
The inference (testing) phase
After the training phase, the selection network $\zeta$ is capable of selecting the most important and phishing-relevant sentences of a given email data $X$ by offering a high value for the corresponding coordinates $\omega_{i}\left(X;\alpha\right)$, meaning that $\omega_{i}\left(X;\alpha\right)$ represents the influence level of the sentence $\boldsymbol{x}_{i}$. We hence can pick out the most relevant sentences based on the magnitude of $\omega_{i}\left(X;\alpha\right)$. Using the selected information, the trained classifier then can predict the vulnerability of the associated email data. In our paper, with the aim of providing a highly qualified and concise explanation of the vulnerability of email data to users, we primarily assess the model’s performance based on the most important (top-1) selected sentence from each email.
Input: An email dataset $S=\left\{\left(X_{1},Y_{1}\right),\dots,\left(X_{N_{S}},Y_{N_{S}}\right)\right\}$ where each email $X_{i}=[\mathbf{x}_{1},...\mathbf{x}_{L}]$ consisting of $L$ sentences while its vulnerability label $Y_{i}\in\left\{0,1\right\}$ (i.e., $1$: phishing and 00: benign).
We denote the number of training iterations $nt$; the mini-batch size $m$; the trade-off hyper-parameter $\lambda$.
We randomly partition $S$ into three different sets including the training set $S_{train}$ (for training the model), the validation set $D_{val}$ (for saving the model during the training process), and the testing set $D_{test}$ (for evaluating the model performance).
1 We initialize the parameters $\alpha$ and $\beta$ of the selection model $\zeta$ (i.e., parameterized by $\omega(.,\alpha)$) and the classifier model $f(.,\beta)$, respectively. for $t=1$ to $nt$ do
2       Choose a mini-batch of embedded phishing-relevant emails denoted by $\{(X_{i},Y_{i})\}_{i=1}^{m}$. Update the classifier’s parameter $\beta$ via minimizing the following cross-entropy loss $\mathcal{L}_{ce}\mathbb{E}_{X,Y}\mathbb{E}_{\mathbf{r}\sim B(0.5)}[\mathcal{L}_{ce}(Y,f_{\beta}(X\odot\mathbf{r})]$ using the Adam optimizer (Kingma and Ba, 2014). Note that minimizing this function is equivalent to maximizing the objective function in Eq. (10). Update the classifier’s parameter $\beta$ and the selection model parameter’s $\alpha$ via minimizing the following objective function $\mathbb{E}_{X,Y}\mathbb{E}_{\mathbf{z}\sim q_{sel}(\mathbf{z}|X;\alpha)}[\mathcal{L}_{ce}(Y,f_{\theta}(X\odot\mathbf{z}))]+\lambda\mathbb{E}_{X}[\sum_{i=1}^{L}\frac{\omega_{i}\left(X;\alpha\right)}{2\sigma^{2}}\left\|\boldsymbol{x}_{i}^{2}\right\|]$ using the Adam optimizer.
3 end for
Output: The trained model for phishing attack localization.
Algorithm 1 The algorithm of our proposed AI2TALE  method for the phishing attack localization problem.
4. Experiments
4.1. Experimental Designs
The fundamental goal of this experiment section is to evaluate our proposed AI2TALE  method and compare it with several potential baseline approaches for phishing attack localization. Below, we present the main research questions of our paper.
(RQ1) Can our AI2TALE  method successfully solve the phishing attack localization in terms of predicting and giving explanations in a weakly supervised setting? Can the selected information (i.e., sentences) also successfully reflect the human cognitive principles used in phishing emails?
Research on phishing detection is one of the most popular solutions for defeating phishing attacks. There have been many machine learning-based and deep learning-based approaches proposed to predict the vulnerability (i.e., phishing or benign) of the associated data (e.g., emails). Although achieving promising performances, AI-based phishing detection methods lack an explanation of the vulnerability prediction (i.e., especially in identifying the specific information causing the data to be classified as phishing). To this end, in this paper, we propose an innovative deep learning-based method for phishing attack localization on email data (the most common phishing way) where we not only aim to detect the vulnerability of the data but also learn and figure out the important and phishing-relevant information existing in the phishing data to provide corresponding comprehensive explanations (i.e., about the key information causing the data phishing). That helps tackle the lack of explainability (transparency) existing in phishing email attack detection. We compare the performance of our AI2TALE  method with several state-of-the-art baselines from explainable machine learning.
Via this research question, we mainly aim to investigate the effectiveness of our AI2TALE  method and baselines in solving the phishing attack localization problem. In particular, we want to determine their success and capability in predicting the vulnerability of email data as well as in learning and identifying the key important and phishing-relevant information (e.g., sentences) that also can reflect the human cognitive principles/triggers (i.e., the definition of ”principles of influence”, namely Reciprocity, Consistency, Social Proof, Authority, Liking, and Scarcity) used in phishing emails (e.g., (Akbar, 2014; Butavicius et al., 2015; Ferreira et al., 2015; Heijden and Allodi, 2019)).
(RQ2) Does the top-$1$ important selected sentences from our AI2TALE  method focus more on the commonly used and important cognitive principles (i.e., Scarcity, Consistency, and Authority) in phishing emails compared to those of the baselines?
As studied and pointed out in phishing attack-relevant research (Akbar, 2014; Butavicius et al., 2015; Ferreira et al., 2015; Heijden and Allodi, 2019), among cognitive triggers used in phishing attacks, some of them (Authority, Scarcity, and Consistency) are more popular than others (Reciprocity, Social, Proof, and Liking). Therefore, in this research question, we investigate how our AI2TALE  method compared to the baselines in focusing and detecting more important and popular cognitive triggers used in phishing emails through the top-$1$ important selected sentences.
(RQ3) Can our proposed information bottleneck theory training term and data-distribution mechanism be successful in improving the data representation learning and supporting the process of learning and figuring out the true phishing-relevant sentences in phishing emails?
The training principle in Eq. (1) induced from the mutual information between the selected information and the corresponding label brings benefits for selecting the important and phishing-relevant sentences from phishing emails, however, as mentioned in Section 3.2.2, it has two potential limitations regarding ”obtaining a superset of phishing-relevant sentences” and ”encoding the vulnerability label via its selections instead of via truly meaningful information”.
To this end, we propose two innovative solutions including the information bottleneck theory training term and data-distribution mechanism. In short, our innovative information bottle-neck training mechanism ensures that only important and label-relevant information (i.e., sentences) will be kept and selected while the data-distribution mechanism ensures the learnable selection process respecting the data distribution to select the meaningful and label-relevant information. It also aids in improving the generalization capability of the selection network, thereby enhancing the performance of the classifier through a random selection process that incorporates data augmentation (please refer to Section 3.2.2 for our detailed explanation).
In this research question, we aim to assess the effectiveness of our proposed information bottleneck theory training term (mentioned in Eq.(9)) and data-distribution mechanism (mentioned in Eq. (10)) in improving the capability of our AI2TALE  method in the process of learning and figuring out the important and true phishing-relevant information (i.e., sentences) in phishing-relevant emails.
4.1.1. Studied datasets
We conducted the experiments on seven phishing-relevant email datasets including IWSPA-AP (i.e., the dataset was collected as part of a shared task to try and address phishing scam emails), Nazario Phishing Corpus (i.e., the dataset of phishing emails (received by one user) that are surrounded by HTML code and need to be stripped to get the useful text data), Miller Smiles Phishing Email (i.e., the dataset contains the bodies of the phishing email scams), Phish Bowl Cornell University (i.e., the dataset contains phishing emails that have been spotted and reported by students and staff at Cornell University), Fraud emails (i.e., the dataset contains fraudulent emails attempting Nigerian Letter where all the emails are in one text file and contain a large amount of header data), Cambridge (i.e., the data set contains a large number of email headers (involving information such as sending and receiving addresses and email subjects) and the body content of phishing emails), and Enron Emails (i.e., the email dataset was released as part of an investigation into Enron, which consists of emails from mostly senior management of Enron).
These datasets contain the common email header and email body. We found that most of this email data is stored in the form of web pages on public websites. We re-processed the datasets before using them for the training and testing processes in terms of removing non-ASCII characters and unrelated information to the contents of the emails such as (e.g., HTML language). We also removed duplicated or similar emails. In the end, we kept 20,000 phishing emails and 20,000 benign emails to conduct our comprehensive and rigorous experiments.
4.1.2. Data processing and embedding
We preprocessed the datasets before injecting them into deep neural networks of our AI2TALE  method and baselines. In the context of our paper, we view each email as a sequence of sentences and we aim to learn and figure out the most important sentence in each email contributing to the vulnerability label (phishing or benign) of the email.
We embedded sentences of each email into vectors. For instance, consider the following sentence ”Your account has been suspended.”, to embed this sentence, we tokenized it to a sequence of tokens including ”your”, ”account”, ”has”, ”been”, ”suspended”, and ”.” using the common Natural Language Toolkit (NLTK). We then used a 150-dimensional token Embedding layer followed by a Dropout layer with a dropped fixed probability $p=0.2$, a 1D convolutional layer with the filter size $150$ and kernel size $3$, and a 1D max pooling layer to encode each sentence. Finally, a mini-batch of emails in which each email consisting of $L$ encoded sentences was fed to our proposed AI2TALE  method and the baselines.
In our paper, the length of each email is padded or truncated with $L=100$ sentences (i.e., we base on the quantile values of the emails’ length of the used datasets to decide the length of each email). We observe that almost all important information relevant to the vulnerability lies in the $100$ first sentences or even lies in some very first sentences.
4.1.3. Measures
To the best of our knowledge, our method is one of the very first approaches for the phishing attack localization problem aiming to tackle and improve the explainability (transparency) of phishing detection. Therefore, studying appropriate measures for phishing attack localization is necessary and also is one of the contributions of our work. To evaluate the performance of our AI2TALE  method and the baselines in phishing attack localization, based on the explainable machine learning and email phishing domain knowledge, we introduce two main metrics to measure the quality of the most important selected sentence in each phishing-relevant email including Label-Accuracy and Cognitive-True-Positive.
Via the Label-Accuracy metric, we measure whether the selected sentences obtained from each model can accurately predict the true vulnerability label (i.e., phishing or benign) of the associated email. The intuition is that the most important sentences contribute the most to the email’s vulnerability, especially to phishing emails. In our experiments, we assess each model’s top-1 selected sentence and measure if this sentence can effectively predict the email’s vulnerability without considering all sentences in the email. The higher the value of the Label-Accuracy measure, the better the model’s performance in figuring out and selecting crucial label-relevant information from the data.
For the Cognitive-True-Positive metric, we investigate if the top-$1$ selected sentence successfully reflects the human cognitive principles used in phishing emails by capturing the cognitive triggers (i.e., Reciprocity, Consistency, Social Proof, Authority, Liking, and Scarcity (Akbar, 2014; Butavicius et al., 2015; Ferreira et al., 2015; Heijden and Allodi, 2019) based on the associated keywords) used in the phishing email. In particular, in our experiments, we consider the top-$1$ selected sentences of phishing emails and calculate how many percent of these selected sentences consist of cognitive triggers.
It is worth noting that to measure if the top-$1$ selected sentence obtained from each method can also reflect and capture the human cognitive principles (triggers) used in each phishing email, we are particularly based on the keywords and phrases often used in each cognitive principle. To obtain the keywords and phrases, for each cognitive principle, we first base on its definition (descriptions) pioneered in the well-known work (Cialdini, 1984) (widely cited and used in phishing-related studies (Akbar, 2014; Butavicius et al., 2015; Ferreira et al., 2015; Heijden and Allodi, 2019)). We then use ChatGPT (OpenAI, 2020) to obtain all the possible keywords and phrases related to each cognitive principle’s definition (descriptions). For example, to the Scarcity principle, some of the related keywords and phrases can be relevant to Time-sensitive language (e.g., ”Act now,” or ”Expires soon”), Limited availability (e.g., ”Only a few left”), and Threats of consequences (e.g., ”Will be deleted” or ”Will lose access”). By relying on the definition (descriptions) of each cognitive principle with the help of ChatGPT in finding all the possible and main keywords and phrases, related to the definition (descriptions) of each cognitive principle, used in the Cognitive-True-Positive measure, we ensure this process is reliable and objective.
Among cognitive triggers used in phishing attacks, some of them (Authority, Scarcity, and Consistency) are more popular than others (Reciprocity, Social, Proof, and Liking). To investigate how our AI2TALE  method compared to the baselines in focusing and detecting more important and popular cognitive triggers. We introduce the SAC measure, standing for the measure related to Scarcity, Authority, and Consistency of cognitive principles. In particular, we consider the top-1 selected sentences of phishing emails and calculate how many percent of these selected sentences successfully consist of the common cognitive triggers.
4.1.4. Baseline methods
The main baselines of our method are state-of-the-art interpretable machine learning approaches for phishing attack localization (i.e., where we not only aim to detect the vulnerability label (i.e., phishing or benign) of phishing-related data but also figure out and highlight the phishing information existing in the data to provide comprehensive explanations about the corresponding phishing vulnerability). It is worth noting that our proposed AI2TALE  method and the used baselines can work on the weakly supervised setting where during the training process, we only use the vulnerability label of the data and do not need to use the ground truth of phishing information in the data for the task of phishing attack localization.
Interpretable machine learning methods are strongly suitable for phishing attack localization in figuring out the most important features of the data obtained from the model to explain the model’s predictive decision. We compared the performance of our AI2TALE  method with several popular and advanced interpretable machine learning approaches including L2X ((Chen et al., 2018b)), INVASE ((Yoon et al., 2019a), ICVH ((Nguyen et al., 2021a), VIBI ((Bang et al., 2021b)), and AIM ((Vo et al., 2023)).
We briefly summarize the baselines as follows:
  • L2X ((Chen et al., 2018b)). An efficient instance-wise feature selection method leverages mutual information for model interpretation. L2X aims to extract a subset of features that are most informative for each given example and the associated model prediction response.
  • INVASE ((Yoon et al., 2019a). Another effective approach for instance-wise feature selection method. INVASE consists of three neural networks including a selector network, a predictor network, and a baseline network which are used to train the selector network using the actor-critic methodology.
  • ICVH ((Nguyen et al., 2021a)). One of the first interpretable deep learning-based methods applied for source code vulnerability localization. ICVH is based on mutual information and a multi-Bernoulli distribution selection process for selecting vulnerability-relevant source code statements.
  • VIBI ((Bang et al., 2021b)). VIBI is a system-agnostic method providing a brief and comprehensive explanation by adopting an information-theoretic principle, the information bottleneck principle, as a criterion for finding such explanations.
  • AIM ((Vo et al., 2023)). A recent innovative additive instance-wise framework for model interpretation. AIM integrates both feature attribution (producing relative importance scores to each feature) and feature selection (directly identifying the subset of features most relevant to the model behavior being explained) into an effective framework for multi-class model interpretation.
4.1.5. Model’s configuration
For the L2X (Chen et al., 2018b), ICVH (Nguyen et al., 2021a), VIBI ((Bang et al., 2021b)), and AIM ((Vo et al., 2023)) methods, they were proposed to explain the output of black-box learning models. To use these methods for phishing attack localization, we keep their principles and train the models to directly approximate $p(Y\mid X)$ using $p(Y\mid\tilde{X})$ where $\tilde{X}$ consists of the selected sentences while $Y$ is ground truth label of the data $X$ instead of the output from the black-box model.
Note that the L2X, ICVH, VIBI, and AIM methods were also proposed to work with sequential text data. Therefore, to these methods, for the architecture of the selection network obtaining $\tilde{X}$ as well as the classifier working on $\tilde{X}$ to mimic $p(Y\mid X)$, we follow the structures used in the corresponding original papers with the same suggested value ranges for hyper-parameters. For the INVASE method, because it was originally designed for working with tabular data, to let it be able to be applicable for sequential text data, we keep its principle and use the selection network as in our AI2TALE  method.
We implemented our AI2TALE  method in Python using Tensorflow (Abadi et al., 2016). The trade-off parameter $\lambda$ is in $\{10^{-1},10^{-2},10^{-3}\}$ while $\sigma$ is in $\{10^{-1},2\times 10^{-1},3\times 10^{-1}\}$. For the networks $\omega\left(\cdot;\alpha\right)$ and $f\left(\cdot;\beta\right)$, we used deep feed-forward neural networks having three and two hidden layers with the size of each hidden layer in $\left\{100,300\right\}$, respectively. The dense hidden layers are followed by a ReLU function as nonlinearity and Dropout (Srivastava et al., 2014) with a retained fixed probability $p=0.8$ as regularization. The last dense layer of the network $\omega\left(\cdot;\alpha\right)$ for learning a discrete distribution is followed by a sigmoid function while the last dense layer of the network $f\left(\cdot;\beta\right)$ is followed by a softmax function for predicting. The temperature $\tau$ for the Gumbel softmax distribution is in $\{0.5,1.0\}$.
For our AI2TALE  method and baselines, we employed the Adam optimizer (Kingma and Ba, 2014) with an initial learning rate of $10^{-3}$, while the mini-batch size is $128$. We split the data set into three random partitions. The first partition contains $80\%$ for training, the second partition contains $10\%$ for validation and the last partition contains $10\%$ for testing. We used $10$ epochs for the training process. We additionally applied gradient clipping regularization to prevent over-fitting. For each method, we ran the corresponding model several times and reported the averaged Label-Accuracy, Cognitive-True-Positive, and SAC measures. We ran our experiments on a 13th Gen Intel(R) Core(TM) i9-13900KF having 24 CPU Cores at 3.00 GHz with 32GB RAM, integrated Gigabyte RTX 4090 Gaming OC 24GB. The source code and data samples for reproducing the experiments of our AI2TALE  method are published at https://anonymous.4open.science/r/AI2TALE/.
4.2. Experimental results
RQ1: Can our AI2TALE  method successfully solve the phishing attack localization in terms of predicting and giving explanations in a weakly supervised setting? Can the selected information (i.e., sentences) also successfully reflect the human cognitive principles used in phishing emails?
Approach
We compare the performance of our AI2TALE  method with the baselines including L2X ((Chen et al., 2018b)), INVASE ((Yoon et al., 2019a), ICVH ((Nguyen et al., 2021a), VIBI ((Bang et al., 2021b)), and AIM ((Vo et al., 2023)) in the task of phishing attack localization (in terms of not only predicting the vulnerability of the email data but also figuring out and highlighting the phishing-relevant information existing in the data. The selected information helps provide comprehensive explanations about the vulnerability of the phishing email data for the users) using the Label-Accuracy and Cognitive-True-Positive measures.
Quantitative Results
The experimental results in Table 1 show that our AI2TALE  method obtains the best performance on both Label-Accuracy and Cognitive-True-Positive compared to the baselines. In particular, our AI2TALE  method performs better than the baselines from nearly (1% to 3%) and (1% to 4%) in Label-Accuracy and Cognitive-True-Positive, respectively. The results demonstrate the effectiveness and advancement of our method for phishing attack localization in learning and figuring out the meaningful and crucial information leading to the vulnerability of the email data, especially for phishing ones, compared to the baselines.
Table 1. The performance of our AI2TALE  method and the baselines for the Label-Accuracy and Cognitive-True-Positive measures on the testing set. We denote the Cognitive-True-Positive as Cognitive-TP for short.
Methods Label-Accuracy Cognitive-TP
INVASE (Yoon et al., 2019a) 98.30% 97.20%
ICVH (Nguyen et al., 2021a) 96.72% 98.10%
L2X (Chen et al., 2018b) 98.25% 97.20%
VIBI (Bang et al., 2021b) 96.65% 94.99%
AIM (Vo et al., 2023) 98.40% 97.10%
AI2TALE  (Ours) 99.33% 98.95%
Table 2. The ground truth and the predicted vulnerability label based on the selected sentence (highlighted in yellow) of each email from our AI2TALE  method.
Truth Model Key words
phishing
phishing
The email content:Someone can access your paypal account, so please confirm your identity to protect your account. Start shopping faster by adding a payment method email customer paypal secure? Now check the account information that belongs to you! Why is my account access limited? Your account access has been limited for the following reason (s) we need to confirm some of your account information. Your case id, for this reason, is pp 009 536 987 252. We face a problem in the ratification of the real owner of the account and for this reason. We require your prompt attention and cooperation.”
Note: This phishing email tries to convey a sense of urgency and asks the email user to confirm their identity and provide account information. It also creates a sense of panic (loss) if the user does not act quickly. Our method successfully figures out the sentence representing the key message of the phishing email for a short and comprehensive explanation.
phishing
phishing
The email content: ”Bulk attention! Your discover account will close soon! Dear member, we have faced some problems with your account, so please update the account. If you do not update will be closed. To update your account, just confirm your information. (it only takes a minute). It’s easy 1. Click the link below to open a secure browser window. 2. Confirm that you’re the owner of the account, and then follow the instructions.”
Note: The message from the selected sentence obtained from our method exhibits cognitive triggers commonly associated with phishing attempts used in the phishing email. In particular, the message implies a sense of urgency (concern) via ”problems with your account while ”dear member” aims to establish a connection with the recipient and imply that the message comes from a trusted source. The phrase ”please update the account” creates a sense of familiarity and consistency.
phishing
phishing
The email content: ”Your account limitation. We regret to inform you that access to your account has been temporarily limited. This has been done due to several failed log in attempts. Case id ax 309 05 66 to restore your account please log in correctly. If you fail to log in correctly your account will be suspended for fraud prevention. You will be able to register again for paypal only after you authenticate your profile. We apologize for the inconvenience. This measure was taken for your protection. Paypal security team.”
Note: The selected sentence conveys a sense of urgency by stating that ”access to your account has been temporarily limited.” This urgency can trigger an emotional response and prompt the recipient to take quick action without thorough consideration. Furthermore, the message’s phrasing, ”we regret to inform you,” and ”access to your account has been temporarily limited,” uses action-oriented language that urges the recipient to take immediate action to address the issue. In short, the model has successfully figured out the important sentence consisting of the cognitive principles used in the phishing email.
In addition, the experimental results on the Cognitive-True-Positive measure of our AI2TALE  method and the baselines shown in Table 1 also indicate that our method and the baselines can successfully localize and figure out the cognitive triggers used in phishing emails via the selected sentences. Compared to the baselines, our proposed AI2TALE  method obtains the best performance on the Cognitive-True-Positive measure.
In summary, in the context of explaining machine learning, particularly when dealing with complex datasets such as email phishing datasets under a weakly supervised setting, the observed improvement of (1% to 3%) for Label-Accuracy and (1% to 4%) for Cognitive-True-Positive, especially within the range approaching 99% or 100%, signifies a substantial improvement.
Qualitative Results
In Table 2, we show some short phishing email samples and the top-1 selected sentences of these emails from our AI2TALE  method as well as the predicted vulnerability labels of the corresponding emails based on the selected sentences. Via these qualitative results, our method demonstrates its effectiveness in learning and figuring out the important and phishing-relevant information (i.e., sentences) from phishing emails. That helps provide useful interpretations for the phishing vulnerability prediction. Via these phishing email samples, it is shown that the cognitive triggers used in these phishing emails, which are mainly related to scarcity (i.e., creates a sense of scarcity and urgency), authority (i.e., creates a sense of trust and legitimacy), and consistency (i.e., creates a sense of familiarity and consistency) principles, are successfully represented via the selected sentences. In addition, we also observe that the most important sentences causing a specific email phishing often appear in some first positions of the email to get the users’ attention.
In conclusion for RQ1: The quantitative experimental results in Table 1 on two main measures (i.e., Label-Accuracy and Cognitive-True-Positive) and the qualitative results in Table 2 show the superiority of our AI2TALE  method in achieving high performances for phishing attack localization on the used real-world email datasets over the baselines. Our method demonstrates its effectiveness in predicting the vulnerability (phishing or benign) of emails and in figuring out the phishing information in phishing emails to give a useful explanation about the vulnerability prediction for users.
Please refer to the appendix for the additional qualitative experiments (Section 4.4) of our AI2TALE  method on other phishing email samples.
RQ2: Does the top-$1$ important selected sentences from our AI2TALE  method focus more on the commonly used and important cognitive principles (i.e., Scarcity, Consistency, and Authority) in phishing emails compared to those of the baselines?
Approach
We compare the performance of our AI2TALE  method with the baselines (i.e., L2X (Chen et al., 2018b), INVASE (Yoon et al., 2019a), ICVH (Nguyen et al., 2021a), VIBI (Bang et al., 2021b), and AIM (Vo et al., 2023)) on the SCA measure where we aim to investigate the capability of our method compared to the baselines in focusing and detecting more important and popular cognitive triggers.
Table 3. The performance of our AI2TALE  method and baselines for the SAC measure on the testing set.
Methods SAC
INVASE (Yoon et al., 2019a) 88.57%
ICVH (Nguyen et al., 2021a) 83.88%
L2X (Chen et al., 2018b) 88.56%
VIBI (Bang et al., 2021b) 82.79%
AIM (Vo et al., 2023) 83.57%
AI2TALE  (Ours) 89.22%
Results
The results in Table 3 show that our AI2TALE  method obtains the best performance on the SAC measure compared to the baselines. In particular, our AI2TALE  method achieves higher performances than the baselines from nearly 0.5% to 6.5%. These results prove the effectiveness of our method in successfully selecting the more popular and important cognitive triggers used in phishing email data through the top-$1$ selected sentences from phishing emails.
In conclusion for RQ2: The experimental results in Table 3 show the advanced capability of our AI2TALE  method in successfully learning and figuring out the commonly used and important cognitive features existing in phishing emails. That helps provide users with crucial information for understanding to defend against phishing attacks.
RQ3: Can our proposed information bottleneck theory training term and data-distribution mechanism be successful in improving the data representation learning and supporting the process of learning and figuring out the true phishing-relevant sentences in phishing emails?
Approach
We compare the performance of our proposed method to itself in the cases of using and without using our proposed information bottleneck theory additional training term mentioned in Eq. (5) and the data-distribution mechanism described in Eq. (10).
We aim to investigate if these additional terms successfully solve the potential limitations of the training principle mentioned in Eq. (1) for further improving the data representation learning that supports the process of learning and figuring out the important and true phishing-relevant information (i.e., sentences) in phishing-relevant emails. Note that in the case without using our introduced information bottleneck theory additional training and data-distribution mechanism terms, we denote the method as AI2TALE-w/oIBT2D.
Results
Table 4. The performance of our method to itself in the cases of using and without using our information bottleneck theory additional training term and the data-distribution mechanism for the Label-Accuracy and Cognitive-True-Positive measures on the testing set. We denote the Cognitive-True-Positive as Cognitive-TP for short.
Methods Label-Accuracy Cognitive-TP SAC
AI2TALE-w/oIBT2D 98.92% 98.40% 88.46%
AI2TALE 99.33% 98.95% 89.22%
The results shown in Table 4 demonstrate the effectiveness of our proposed information bottleneck theory additional training and the data-distribution mechanism in boosting the data representation learning that facilitates and supports the process of learning and figuring out the important and true phishing-relevant information (i.e., sentences) in phishing-relevant emails.
In particular, our method consists of our introduced bottleneck theory additional training term and the data-distribution mechanism (AI2TALE) obtained better performance (from around 0.5% to 1.0%) than the case without using these terms (AI2TALE-w/oIBT2D) on Label-Accuracy, Cognitive-True-Positive, and SAC measures. In the realm of explaining machine learning when dealing with complex datasets, such as email phishing datasets, under weakly supervised conditions, these improvements are promising. That exhibits the positive effect of the bottleneck theory additional training and data-distribution mechanism terms in improving the overall performance of our AI2TALE  method to obtain state-of-the-art performance on the Label-Accuracy, Cognitive-True-Positive, and SAC measures compared to the baselines.
In conclusion for RQ3: The experimental results in Table 4 show the effectiveness of our bottleneck theory training term and the data-distribution mechanism in boosting the data representation learning process that is beneficial for not only the emails’ vulnerability prediction but also for automatically learning and figuring out the truly crucial phishing-relevant information in phishing emails compared to the case without using these terms.
Training time
The training time of our framework in the cases of using and without using our proposed information bottleneck theory training term and data-distribution mechanism is considerably short, costing approximately 36.47 seconds and 21.99 seconds, respectively, for each epoch (as mentioned in Section 4.1.5, we ran our experiments on a 13th Gen Intel(R) Core(TM) i9-13900KF having 24 CPU Cores at 3.00 GHz with 32GB RAM, integrated Gigabyte RTX 4090 Gaming OC 24GB).
4.3. Human evaluation
We additionally conduct a human evaluation to investigate the usefulness of our proposed AI2TALE  method in figuring out important and phishing-related information (e.g., sentences) in phishing emails to help provide comprehensive explanations about the phishing of the corresponding email data. In particular, we evaluate whether the top-1 selected sentence in each phishing email by our proposed AI2TALE  method is perceived as convincing information for email users. To do that, we asked participants to evaluate the selected sentences of $10$ different phishing emails (randomly chosen from the testing set) in terms of whether the sentence selected in each email is important to influence and persuade users to follow the instructions in the email.
Figure 2. Human evaluation on the importance of the top-$1$ selected information (i.e., a sentence) from each email (by our AI2TALE  method) in affecting and persuading users to follow the instructions from the email. We evaluate the selected sentences of 10 different phishing emails (randomly chosen from the testing set).
There were $25$ university students and staff participating in our survey. All participants responded that they have used emails for work and study, and they have both experienced and heard about phishing attacks. Based on the responses, in summary, as depicted in Figure 2, 81% of participants selected either ”Agree” (55%) or ”Strongly Agree” (26%) when asked if they believe the selected sentences affect users’ decision to follow the instructions in the survey phishing emails (note that in each phishing email, we use the top-1 selected sentence obtain from our AI2TALE  method). In contrast, 10% of participants chose ”Neutral”, while 9% chose ”Disagree” or ”Strongly Disagree”. These human evaluations demonstrate the effectiveness of our method in successfully figuring out important phishing information (the attackers use it to deceitfully persuade the target users). The selected information in each phishing helps provide a useful and comprehensive explanation of the vulnerability prediction of the email, especially for the phishing one, to users. In essence, this approach enhances the explainability (transparency) of phishing detection, addressing and improving the overall clarity of the prediction process.
Note that, in this human evaluation, we implemented careful study design protocols to minimize potential priming. To ensure objectivity in the results, no information was provided about the source of the selected sentences.
4.4. Additional quantitative results
In our study, as presented in RQ1, via the Label-Accuracy metric, we aim to measure whether the selected sentences obtained from each model can accurately predict the true vulnerability label (i.e., phishing or benign) of the associated email. The intuition is that the most important sentences contribute the most to the email’s vulnerability, especially to phishing emails. In our experiments, we assess each model’s top-1 selected sentence and measure if it can effectively predict the email’s vulnerability without considering all sentences in the email. The higher the value of the Label-Accuracy measure, the better the model’s performance in selecting crucial label-relevant information from the data.
The datasets used in our paper are balanced, with an equal number of phishing and benign emails (refer to the Studied datasets section 4.1.1 for details). In the context of a balanced dataset, the accuracy measure is sufficient, and high accuracy implies low false negative (FN) and false positive (FP). To fulfill our experiments on this aspect, we calculated the F1-score for our AI2TALE  method and baselines. Specifically, we computed the F1-score in two scenarios: (i) exclusively on the testing phishing emails, and (ii) on both the testing phishing and benign emails. The results for our AI2TALE  method and the baselines (AIM, VIBI, L2X, ICVH, INVASE) in scenarios (i) and (ii) are shown in Table 5.
Table 5. The performance of our AI2TALE  method and the baselines for the F1-score measure on the testing set in two scenarios (i) and (ii).
Methods F1-score (i) F1-score (ii)
INVASE (Yoon et al., 2019a) 98.313% 98.299%
ICVH (Nguyen et al., 2021a) 96.732% 96.725%
L2X (Chen et al., 2018b) 98.261% 98.249%
VIBI (Bang et al., 2021b) 96.626%96.649%
AIM (Vo et al., 2023) 98.406% 98.399%
AI2TALE  (Ours) 99.324% 99.325%
In summary, the F1-score results for our AI2TALE  method and baselines closely align with those obtained using the accuracy measure. Notably, our AI2TALE  method exhibits consistent effectiveness and advancement, surpassing the baselines by approximately 1% to 3% in both scenarios (i) and (ii).
5. Threats to validity
Construct validity
Key construct validity threats are if our assessments of the proposed method and baselines demonstrate the ability for phishing attack localization. In our paper, we study an important problem of phishing attack localization where we not only aim to automatically learn and figure out the important phishing-relevant information (e.g., sentences) in each phishing email but also can detect the vulnerability $Y$ (i.e., phishing or benign) of the corresponding email based on the crucial selected information. The selected phishing-relevant information (e.g., sentences) helps provide comprehensive explanations about the vulnerability of the phishing email data. To evaluate the performance of our method and baselines, we introduce three main measures including Label-Accuracy, Cognitive-True-Positive, and SAC.
Internal validity
Key internal validity threats are relevant to the choice of hyper-parameter settings (i.e., optimizer, learning rate, number of layers in deep neural networks, etc.). It is worth noting that finding a set of optimal hyperparameter settings of deep neural networks is expensive due to a large number of trainable parameters. To train our method, we only use the common and default values of hyper-parameters such as using Adam optimizer and the learning rate equals $10^{-3}$. We also report the hyperparameter settings in our released reproducible source code samples to support future replication studies.
External validity
Key external validity threats include whether our proposed AI2TALE  method can generalize well to different phishing vulnerabilities and other different types of phishing email datasets. We mitigated this problem by conducting our experiments on seven phishing-relevant email datasets including IWSPA-AP, Nazario Phishing Corpus, Miller Smiles Phishing Email, Phish Bowl Cornell University, Fraud emails, Cambridge, and Enron Emails.
6. Future work
While our study marks a pioneering effort and significantly contributes to advancing transparency and detection capabilities for phishing email attacks, we acknowledge the existence of one challenging case that poses a challenge to our AI2TALE  approach and other phishing detection and localization methods.
In particular, in the broad landscape of email phishing attacks, an exceptionally challenging scenario arises when phishing emails mimic the content of benign emails, differing only in the embedded URLs. For instance, attackers may send phishing emails to victims using identical text content as genuine communications from entities like PayPal, with the sole modification being the replaced URL. In this case, classifying and explaining phishing attacks primarily relying on email content and text within embedded URLs may become particularly challenging. To the best of our knowledge, this problem may need a more comprehensive approach incorporating URL analysis and additional relevant features. This augmentation aims to enhance the overall effectiveness of phishing localization models, ensuring more comprehensive and accurate detection and explanation of phishing attacks. Addressing this issue will be a focal point of our future work.
7. Related background
In this section, we briefly present the main related background used in our proposed AI2TALE  method.
7.1. Mutual information
Mutual information (MI) is used to measure the dependence between two random variables (Shannon, 1998; Cover and Thomas, 2006). It captures how much the knowledge of one random variable reduces the uncertainty of the other. In particular, MI quantifies the amount of information obtained about one random variable by observing the other random variable. For instance, suppose variable A signifies the outcome of rolling a standard 6-sided die, and variable B represents whether the roll yields an even (0 for even, 1 for odd) result. It is evident that information from B offers insights into the value of A, and vice versa. In essence, these random variables exhibit mutual information.
Assume that we have two random variables $X$ and $Y$ drawn from the joint distribution $p(x,y)$ with two corresponding marginal distributions $p(x)$ and $p(y)$. The mutual information between $X$ and $Y$ denoted by $I(X,Y)$ is the relative entropy between the joint distribution $p(x,y)$ and the product distribution $p(x)p(y)$, and is defined as follows:
$I(X,Y)$ $=\int p(x,y)\log\frac{p(x,y)}{p(x)p(y)}dxdy$
$=D_{KL}(p(x,y)||p(x)p(y))$
where $D_{KL}(p(x,y)||p(x)p(y))$ is the Kullback-Leibler divergence measuring the relative entropy (i..e, the difference in information) represented by two distributions, i.e., the product of marginal distributions $p(x)p(y)$ of $X$ and $Y$ from their joint distribution $p(x,y)$.
7.2. Information bottleneck theory
Here, we consider the supervised learning context where we want to predict corresponding outputs (e.g., labels) $\left\{{\mathbf{y}}_{i}\right\}_{i=1}^{n}$ of given inputs $\left\{{\mathbf{x}}_{i}\right\}_{i=1}^{n}$. A deep learning network (DNN) will learn latent representations (i.e., latent features in the latent space that contain useful information to describe the data) $\left\{\tilde{{\mathbf{x}}}_{i}\right\}_{i=1}^{n}$ of the corresponding input data samples $\left\{{\mathbf{x}}_{i}\right\}_{i=1}^{n}$ in terms of enabling good predictions and generalizations.
Assume that the whole hidden layer in Figure 3 is denoted by a random variable $\tilde{X}$ while the input and output layers are denoted by random variables $X$ and $Y$ respectively. We can describe this hidden layer by two conditional distributions: the encoder $p(\tilde{{\mathbf{x}}}|{\mathbf{x}})$ and the decoder $p({\mathbf{y}}|\tilde{{\mathbf{x}}})$. This transformation process preserves the information of the input layer $X$ without considering which individual neurons within the hidden layer $\tilde{X}$ encode which features (i.e., neurons) of $X$. An optimal encoder process of the mutual information between $X$ and the desired output $Y$ denoted by $I(X,Y)$ can create the most compact encoding (i.e., minimally sufficient statistic) $\tilde{X}$ of the input data $X$ while $\tilde{X}$ still has enough information (i.e., $\tilde{X}$ can capture the important features of $X$ as well as remove the unnecessary parts of $X$ that do not make any contributions to the prediction of $Y$) to predict $Y$ as accurately as possible.
Figure 3. An architecture of a simple deep neural network in a supervised learning context for the classification problem.
An information bottleneck (Tishby et al., 2000; Naftali Tishby, 2015) is proposed to be a computational framework that aims to find the most compact encoding $\tilde{X}$ of the input data $X$. In particular, it is the optimal trade-off between the compression $\tilde{X}$ and the prediction of the desired output $Y$ as described in the following optimization problem:
$\min_{p(\tilde{{\mathbf{x}}}|{\mathbf{x}}),p({\mathbf{y}}|\tilde{{\mathbf{x}}}),p(\tilde{{\mathbf{x}}})}\left\{I(X,\tilde{X})-\beta I(\tilde{X},Y)\right\}$
where $\beta$ specifies the amount of relevant information captured by the encoding process (i.e., the representations $\tilde{X}$ and $I(\tilde{X},Y)$).
8. Conclusion
In this paper, we have successfully proposed one of the very first innovative deep learning-based methods derived from an information-theoretic perspective and information bottleneck theory for the phishing attack localization problem where automated AI-based techniques have not yet been well studied. Our AI2TALE  method can work effectively in the weakly supervised setting providing an important practical solution to solve the problem in terms of not only accurately predicting the vulnerability of the email data but also automatically learning and figuring out the important and phishing-relevant information in each phishing email. The selected information provides useful explanations about the vulnerability of the corresponding phishing email data. In addition, we also introduce appropriate measures for phishing attack localization. The rigorous and comprehensive experiments on seven real-world email datasets show the superiority of our proposed AI2TALE  method over the state-of-the-art baselines.